Circular No. 6580

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 T. Boller and W. Voges, Max-Planck-Institut für Extraterrestrische

Physik, Garching; F. Frontera, Universitá di Ferrara and Instituto Tecnologie e Studio delle Radiazioni Extraterrestri, CNR, Bologna; and E. Costa and L. Piro, Istituto Astrofisica Spaziale, Frascati, report: “Within the 3′-radius error box of the BeppoSAX/WFC position for this γ-ray burst (IAUC 6572), no source is detected in the ROSAT all-sky-survey data between 1990 Aug. 26 and 28. The survey exposure time at the BeppoSAX/WFC position is 435 s. The ROSAT 1.7σ upper limit in this error box is 0.0272 count/s (0.1–2.4 keV), which corresponds to 9.2 × 10^{−13} erg cm^{−2} s^{−1}, assuming a column density of hydrogen N_{H} = 1.6 × 10^{21} cm^{−2} and a power-law model with energy index 0.8. The distance to the nearest ROSAT all-sky-survey source from the GRB centroid is 12′.7. Within the 50′′ error box of the BeppoSAX Narrow Field Instruments centered on α = 5^{h}01^{m}44^{s}, δ = 11^{o}46′.7 (IAUC 6576), the ROSAT all-sky-survey 2.5σ upper limit is 0.0056 count/s (0.1–2.4 keV), which corresponds to 1.9 × 10^{−13} erg cm^{−2} s^{−1}; this is compatible with the BeppoSAX flux level reported for GRB 970228 on Mar. 3.734 UT (IAUC 6576).”

H. Pedersen, Copenhagen University Observatory; S. Benetti and M. Turatto, European Southern Observatory (ESO); P. Goudfrooij and A. Marconi, Space Telescope Science Institute, Baltimore; P. Jablonka, Département d’Astrophysique Extragalactique et de Cosmologie, Observatoire de Paris; E. Costa, Istituto Astrofisica Spaziale, CNR, Frascati; F. Frontera, Universitá di Ferrara; E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; J. Heise, Space Research Organization of the Netherlands, Utrecht; J. Studt, Hamburg Observatory; and A. J. Castro-Tirado, LAEFF, Madrid, report: “We have observed the position of the VLA source (IAUC 6576) using the ESO 3.6-m telescope (+ EFOSC1 instrument). On the sharpest and deepest exposure (R, 3 min, Mar. 8), the optical object reported by Margon et al. (IAUC 6577) has the appearance of a centrally condensed galaxy. Integrated within a 5′′ diaphragm, we find B > 22, R = 21.06 (Mar. 8); and V = 22.27, R = 21.05 (Mar. 9). The measurement errors are 0.1–0.2 mag. Several fainter galaxies are close by; one is 2′′.5 east, another 5′′ west. Low-dispersion, noisy spectra (range 380– 700 and 600–1000 nm) obtained on Mar. 9 show no evident emission or absorption lines present in the red continuum.”

1997 March 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green